# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v2

@@ 1.3. Results toward the non-abelian Cohen–Lenstra heuristics
Let $K$ be a quadratic extension of $\mathbb{F}_{q}(t)$, let $K^{\sharp}_{\emptyset}/K$ denote the maximal unramified extension of $K$ that is split completely over $\infty$, in the sense of § 2.3, and has degree prime to $2q$. Define $G^{\sharp}_{\emptyset}(K):=\operatorname{Gal}(K^{\sharp}_{\emptyset}/K)$. Then, $G^{\sharp}_{\emptyset}(K)$ has an action of $\mathbb{Z}/2\mathbb{Z}$ coming from the $\mathbb{Z}/2\mathbb{Z}=\operatorname{Gal}(K/\mathbb{F}_{q}(t))$ action on $K$. This action is well-defined using the Schur–Zassenhaus lemma; for further explanation, see the paragraph prior to [LWZB24, Definition 2.1]. For $H$ a group with a $\mathbb{Z}/2\mathbb{Z}$ action, we use $\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}(G^{\sharp}_{\emptyset}(K),H)$ to denote the set of $\mathbb{Z}/2\mathbb{Z}$ equivariant surjections $G^{\sharp}_{\emptyset}(K)\to H$. We use ^$\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}$ to denote the pro-prime to $2q$ completion of for ^$\widehat{\mathbb{Z}}(1):=\lim_{n}\mu_{n}(\overline{\mathbb{F}}_{q})$. For $L/K$ an extension of fields, there is a certain map $\omega_{L/K}:\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}\to H_{2}(\operatorname{Gal}(L/\mathbb{F}_{q}(t)),\mathbb{Z})_{(2q)^{\prime}}$ defined in [Liu22, Definition 2.13]. For $\pi\in\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}(G^{\sharp}_{\emptyset}(K),H)$, we use $\pi_{*}:H_{2}(\operatorname{Gal}(K^{\sharp}_{\emptyset}/\mathbb{F}_{q}(t)),\mathbb{Z})_{(2q)^{\prime}}\to H_{2}(H\rtimes\mathbb{Z}/2\mathbb{Z},\mathbb{Z})_{(2q)^{\prime}}$ to denote the corresponding map induced by $\pi$. We say that $H$ is admissible for the action of $\mathbb{Z}/2\mathbb{Z}$ if $H$ has odd order and is generated by elements of the form $h^{-1}\cdot\iota(h)$ for $h\in H$ and $\iota$ the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$.
@@
where $[H:H^{\mathbb{Z}/2\mathbb{Z}}]$ denotes the index of the $\mathbb{Z}/2\mathbb{Z}$ invariants of $H$ in $H$.
@@ 1.4.4. In algebraic geometry as a moduli space
Theorem 1.4.5 is known to hold for certain special components $Z$ by work of Bianchi and Miller [BM23, Corollary C’] and the text following it. Via personal communication, we learned this argument was also known to Ellenberg–Venkatesh–Westerland. In particular, this applies when $G$ is the dihedral group of order $2n$, for $n$ odd, to components $Z$ whose boundary monodromy (see § 2.2) is a generator of the index $2$ abelian subgroup of $G$. However, it does not apply, for example, when $G$ is the dihedral group of order $2n$ and the boundary monodromy of $Z$ is trivial. See also [Tie16, §6] for additional computations of the homology of Hurwitz spaces in other special cases similar to the above, via essentially the same strategy.
Additionally, we mention that the results of [BM23, Corollary C’] do not require a non-splitting hypothesis on $(G,c)$. As a trade off, without the non-splitting hypothesis, they obtain bounds on the stability range which are much worse than linear in the homological degree; see also [BM23, §1.2].
@@ 1.5. The stable homology for $S_{3}$ and degree $3$ covers
$h^{i}([\operatorname{CHur}^{S_{3},\{(12),(13),(23)\}}_{n}/S_{3}];\mathbb{Q})=\begin{cases}1&\text{ if }i\in\{0,1\}\\ 0&\text{ otherwise. }\\ \end{cases}$
@@ 1.6. Conjectures on stable homology of Hurwitz spaces
As pointed out to us by Dan Petersen, the rational Chow group of $\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ is 00 in positive degrees and $\mathbb{Q}$ in degree 00, so if § 1.6 were true, this would give an especially simple description of the stable Chow groups of Hurwitz spaces.
@@ 1.9. Acknowledgements
We thank Jordan Ellenberg and Craig Westerland for sharing a number of the ideas they tried while working on this problem. The ideas in paper build heavily upon ideas in the retracted preprint [EVW12], and we were able to draw much inspiration from the ideas presented there. We thank Melanie Wood for suggesting § 1.6, listening to a detailed explanation of the argument, and for helpful comments on the paper. We thank Andrea Bianchi for an extremely helpful suggestion which allowed us to remove an extraneous condition we had on Theorem 4.2.1. We thank Yuan Liu for explaining to us the correct version of the Cohen–Lenstra heuristics in the function field case, and for help understanding the non-abelian Cohen–Lenstra heuristics with roots of unity. We also thank Samir Canning, Kevin Chang, Ben Knudsen, Jef Laga, Jeremy Miller, Sun Woo Park, Anand Patel, Sam Payne, Dan Petersen, Eric Rains, Oscar Randal-Williams, Will Sawin, Mark Shusterman, Agata Smoktunowicz, Nathalie Wahl, and Angelina Zheng for helpful conversations. Landesman was supported by the National Science Foundation under Award No. DMS 2102955. Levy was supported by the NSF Graduate Research Fellowship under Grant No. 1745302, and by the Clay Research Fellowship.
@@ 2.1. The definition of Hurwitz spaces
We begin by carefully defining the Hurwitz spaces we will work with. We start by defining the Hurwitz stack $[\operatorname{Hur}^{G,c}_{n,B}/G]$. This will be a quotient by an action of $G$ of the pointed Hurwitz space $\operatorname{Hur}^{G,c}_{n,B}$, which we will define next, and this explains why there is a quotient by $G$ in the notation for $[\operatorname{Hur}^{G,c}_{n,B}/G]$.
@@
Let $G$ be a finite group, $c\subset G$ be a union of conjugacy classes. Let $B$ be a scheme with $|G|$ invertible on $B$. Assume either that $c$ is closed under invertible powering ($g\in c\implies g^{t}\in c$ for any $t$ prime to $|G|$,) or that $B$ is henselian with residue field $\operatorname{Spec}\mathbb{F}_{q}$ and $c$ is closed under under $q$th powering ($g\in c\implies g^{q}\in c$). Let $[\operatorname{Hur}^{G,c}_{n,B}/G]$ denote the Hurwitz stack whose $T$ points $[\operatorname{Hur}^{G,c}_{n,B}/G](T)$ is the groupoid
@@
    Two such covers are considered equivalent if they are related by the $G$-conjugation action.
  7. (7)
    The morphisms between two points $(D_{i},i_{i},X_{i},h_{i})$ for $i\in\{1,2\}$ are given by $(\phi_{D},\psi_{X})$ where $\phi_{D}:D_{1}\simeq D_{2}$ is an isomorphism so that $i_{2}\circ\phi_{D}=i_{2}$ and $\psi_{X}:X_{1}\simeq X_{2}$ is an isomorphism such $h_{2}\circ\psi_{X}=h_{1}$ and $\psi_{X}=g^{-1}\psi_{X}g$ for every $g\in G$.
@@
The above paragraph shows that $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ is an algebraic stack, but does not yet justify the existence of the stacks $[\operatorname{Hur}^{G,c}_{n,B}/G]$ under the suitable powering hypotheses, depending on $B$. First, if $|G|$ is invertible on $B$, and $c$ is closed under invertible powering, then $[\operatorname{Hur}^{G,c}_{n,B}/G]$ is a union of components of the Hurwitz space $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ by [LWZB24, Corollary 12.2]. Second, we explain why $[\operatorname{Hur}^{G,c}_{n,B}/G]$ is a union of components of the Hurwitz space $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ when $B$ is a Henselian dvr with finite residue field. We can first reduce to the case that $B=\mathbb{F}_{q}$ using that $B$ is Henselian and $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ has a smooth proper compactification by [EL23, Corollary B.1.4]. Since $c$ is closed under $q$th powering, the result then holds using by [LWZB24, Theorem 12.1(2)].
@@
Fix an integer $w$ and define $\mathscr{P}^{w}$ to be the root stack of order $w$ along $\infty$ of $\mathbb{P}^{1}$. (See [Cad07, Definition 2.2.4], where this is notated as (,$\mathbb{P}^{1}_{(\infty,w)}$.) The fiber of this root stack over $\infty$ is the stack quotient $[\left(\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})\right)/\mu_{r}]$ of the relative spectrum $\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})$ by $\mu_{r}$. Let ~$\widetilde{\infty}:B\to\mathscr{P}^{w}$ denote the section over $\sigma$ corresponding to map $B\to[\left(\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})\right)/\mu_{r}]$ given by the trivial $\mu_{r}$ torsor over $B$, $\mu_{r}\to B$, and the $\mu_{r}$ equivariant map $\mu_{r}\to B\to\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})$. We use notation as in § 2.1,
@@
where $D,i,X,$ and $h$ satisfy the properties listed in § 2.1. We also assume the order of inertia of $h$ along $\infty$ is $w$ and define ~$\widetilde{\infty}_{T}$ to be the base change of the section $\widetilde{\infty}$ defined above to $T$. We additionally impose the condition that $h^{\prime}$ is a finite locally free $G$-cover, étale over ~$\widetilde{\infty}_{T}$, such that the composition of $h^{\prime}:X\to\mathscr{P}^{w}_{T}$ with the coarse space map $\mathscr{P}^{w}_{T}\to\mathbb{P}^{1}_{T}$ is $h$, and $t:T\to X\times_{h^{\prime},\mathscr{P}^{w}_{T},\widetilde{\infty}_{T}}T$ is a section of $h^{\prime}$ over ~$\widetilde{\infty}_{T}$.
@@
In the definition of pointed Hurwitz space, § 2.1, we choose a particular section ~$\widetilde{\infty}:B\to\mathscr{P}^{w}$ over $\infty:B\to\mathscr{P}^{1}_{B}$. However, one can show that if we chose a different section ~$\widetilde{\infty}^{\prime}:B\to\mathscr{P}^{w}$, the resulting pointed Hurwitz space would be isomorphic. The idea is that the sections $\widetilde{\infty}$ and $\widetilde{\infty}^{\prime}$ factor respectively through maps $\alpha:B\to[B/\mu_{w}]$ and $\alpha^{\prime}:B\to[B/\mu_{w}]$, and so it suffices to produce a map $\beta:[B/\mu_{w}]\to[B/\mu_{w}]$ so that $\alpha\circ\beta=\alpha^{\prime}$, as this map will induce a map taking $\widetilde{\infty}$ to $\widetilde{\infty}^{\prime}$, which then induces an isomorphism of these two Hurwitz spaces. The desired map $\beta$ is simply given as the composition of the structure map $[B/\mu_{w}]\to B$ with $\alpha^{\prime}$.
@@
We use $\operatorname{CHur}^{G,c}_{n,B}$ to denote the open and closed subscheme of $\operatorname{Hur}^{G,c}_{n,B}$ parameterizing covers $X\to\mathbb{P}^{1}_{T}$ so that the geometric fibers of $X$ over $T$ are connected. We use $[\operatorname{CHur}^{G,c}_{n,B}/G]$ to denote the open and closed substack of $[\operatorname{Hur}^{G,c}_{n,B}/G]$ parameterizing covers $X\to\mathbb{P}^{1}_{T}$ so that the geometric fibers of $X$ over $T$ are connected.
@@
We will often work over the complex numbers, and hence it will be convenient to assume $B=\operatorname{Spec}\mathbb{C}$. Therefore, we define $[\operatorname{Hur}^{G,c}_{n}/G]:=[\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}/G],\operatorname{Hur}^{G,c}_{n}:=\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}},[\operatorname{CHur}^{G,c}_{n}/G]:=[\operatorname{CHur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}/G],\operatorname{CHur}^{G,c}_{n}:=\operatorname{CHur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}$. When $R$ is a ring, we often use use $\operatorname{Hur}^{G,c}_{n,R}$ in place of $\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}R},$ and use analogous notation for other variants of this Hurwitz space.

### v3

@@ 1.3. Results toward the non-abelian Cohen–Lenstra heuristics
Let $K$ be a quadratic extension of $\mathbb{F}_{q}(t)$, let $K^{\sharp}_{\emptyset}/K$ denote the maximal unramified extension of $K$ that is split completely over $\infty$, in the sense of § 2.3, and has degree prime to $2q$. Define $G^{\sharp}_{\emptyset}(K):=\operatorname{Gal}(K^{\sharp}_{\emptyset}/K)$. Then, $G^{\sharp}_{\emptyset}(K)$ has an action of $\mathbb{Z}/2\mathbb{Z}$ coming from the $\mathbb{Z}/2\mathbb{Z}=\operatorname{Gal}(K/\mathbb{F}_{q}(t))$ action on $K$. This action is well-defined using the Schur–Zassenhaus lemma; for further explanation, see the paragraph prior to [LWZB24, Definition 2.1]. For $H$ a group with a $\mathbb{Z}/2\mathbb{Z}$ action, we use $\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}(G^{\sharp}_{\emptyset}(K),H)$ to denote the set of $\mathbb{Z}/2\mathbb{Z}$ equivariant surjections $G^{\sharp}_{\emptyset}(K)\to H$. We use $\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}$ to denote the pro-prime to $2q$ completion of for $\widehat{\mathbb{Z}}(1):=\lim_{n}\mu_{n}(\overline{\mathbb{F}}_{q})$. For $L/K$ an extension of fields, there is a certain map $\omega_{L/K}:\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}\to H_{2}(\operatorname{Gal}(L/\mathbb{F}_{q}(t)),\mathbb{Z})_{(2q)^{\prime}}$ defined in [Liu22, Definition 2.13]. For $\pi\in\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}(G^{\sharp}_{\emptyset}(K),H)$, we use $\pi_{*}:H_{2}(\operatorname{Gal}(K^{\sharp}_{\emptyset}/\mathbb{F}_{q}(t)),\mathbb{Z})_{(2q)^{\prime}}\to H_{2}(H\rtimes\mathbb{Z}/2\mathbb{Z},\mathbb{Z})_{(2q)^{\prime}}$ to denote the corresponding map induced by $\pi$. We say that $H$ is admissible for the action of $\mathbb{Z}/2\mathbb{Z}$ if $H$ has odd order and is generated by elements of the form $h^{-1}\cdot\iota(h)$ for $h\in H$ and $\iota$ the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$.
@@
where $H:H^{\mathbb{Z}/2\mathbb{Z}}]$ denotes the index of the $\mathbb{Z}/2\mathbb{Z}$ invariants of $H$ in $H$.
@@ 1.4.4. In algebraic geometry as a moduli space
Theorem 1.4.5 is known to hold for certain special components $Z$ by work of Bianchi and Miller [BM25, Corollary C’] and the text following it. Via personal communication, we learned this argument was also known to Ellenberg–Venkatesh–Westerland. In particular, this applies when $G$ is the dihedral group of order $2n$, for $n$ odd, to components $Z$ whose boundary monodromy (see § 2.2) is a generator of the index $2$ abelian subgroup of $G$. However, it does not apply, for example, when $G$ is the dihedral group of order $2n$ and the boundary monodromy of $Z$ is trivial. See also [Tie16, §6] for additional computations of the homology of Hurwitz spaces in other special cases similar to the above, via essentially the same strategy.
Additionally, we mention that the results of [BM25, Corollary C’] do not require a non-splitting hypothesis on $(G,c)$. As a trade off, without the non-splitting hypothesis, they obtain bounds on the stability range which are much worse than linear in the homological degree; see also [BM25, §1.2].
@@ 1.5. The stable homology for $S_{3}$ and degree $3$ covers
[…]
@@ 1.6. Conjectures on stable homology of Hurwitz spaces
As pointed out to us by Dan Petersen, the rational Chow group of $\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ is 0 in positive degrees and $\mathbb{Q}$ in degree 0, so if § 1.6 were true, this would give an especially simple description of the stable Chow groups of Hurwitz spaces.
@@ 1.9. Acknowledgements
We thank Jordan Ellenberg and Craig Westerland for sharing a number of the ideas they tried while working on this problem. The ideas in paper build heavily upon ideas in the retracted preprint [EVW12], and we were able to draw much inspiration from the ideas presented there. We thank Melanie Wood for suggesting § 1.6, listening to a detailed explanation of the argument, and for helpful comments on the paper. We thank Andrea Bianchi for an extremely helpful suggestion which allowed us to remove an extraneous condition we had on Theorem 4.2.1. We thank Yuan Liu for explaining to us the correct version of the Cohen–Lenstra heuristics in the function field case, and for help understanding the non-abelian Cohen–Lenstra heuristics with roots of unity. We also thank Samir Canning, Kevin Chang, Elia Gorokhovsky, Ben Knudsen, Jef Laga, Jeremy Miller, Sun Woo Park, Anand Patel, Sam Payne, Dan Petersen, Eric Rains, Oscar Randal-Williams, Will Sawin, Mark Shusterman, Agata Smoktunowicz, Nathalie Wahl, and Angelina Zheng for helpful conversations. Landesman was supported by the National Science Foundation under Award No. DMS 2102955. Levy was supported by the NSF Graduate Research Fellowship under Grant No. 1745302, and by the Clay Research Fellowship.
@@ 2.1. The definition of Hurwitz spaces
We begin by carefully defining the Hurwitz spaces we will work with. We start by defining the Hurwitz stack $\operatorname{Hur}^{G,c}_{n,B}/G]$. This will be a quotient by an action of $G$ of the pointed Hurwitz space $\operatorname{Hur}^{G,c}_{n,B}$, which we will define next, and this explains why there is a quotient by $G$ in the notation for $\operatorname{Hur}^{G,c}_{n,B}/G]$.
@@
Let $G$ be a finite group, $c\subset G$ be a union of conjugacy classes. Let $B$ be a scheme with $|G|$ invertible on $B$. Assume either that $c$ is closed under invertible powering ($g\in c\implies g^{t}\in c$ for any $t$ prime to $|G|$,) or that $B$ is henselian with residue field $\operatorname{Spec}\mathbb{F}_{q}$ and $c$ is closed under under $q$th powering ($g\in c\implies g^{q}\in c$). Let $\operatorname{Hur}^{G,c}_{n,B}/G]$ denote the Hurwitz stack whose $T$ points $\operatorname{Hur}^{G,c}_{n,B}/G](T)$ is the groupoid
@@
    The morphisms between two points $(D_{i},i_{i},X_{i},h_{i})$ for $i\in\{1,2\}$ are given by $(\phi_{D},\psi_{X})$ where $\phi_{D}:D_{1}\simeq D_{2}$ is an isomorphism so that $i_{2}\circ\phi_{D}=i_{1}$ and $\psi_{X}:X_{1}\simeq X_{2}$ is an isomorphism such $h_{2}\circ\psi_{X}=h_{1}$ and $\psi_{X}=g^{-1}\psi_{X}g$ for every $g\in G$.
@@
The above paragraph shows that $\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ is an algebraic stack, but does not yet justify the existence of the stacks $\operatorname{Hur}^{G,c}_{n,B}/G]$ under the suitable powering hypotheses, depending on $B$. First, if $|G|$ is invertible on $B$, and $c$ is closed under invertible powering, then $\operatorname{Hur}^{G,c}_{n,B}/G]$ is a union of components of the Hurwitz space $\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ by [LWZB24, Corollary 12.2]. Second, we explain why $\operatorname{Hur}^{G,c}_{n,B}/G]$ is a union of components of the Hurwitz space $\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ when $B$ is a Henselian dvr with finite residue field. We can first reduce to the case that $B=\mathbb{F}_{q}$ using that $B$ is Henselian and $\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ has a smooth proper compactification by [EL23, Corollary B.1.4]. Since $c$ is closed under $q$th powering, the result then holds using by [LWZB24, Theorem 12.1(2)].
@@
Fix an integer $w$ and define $\mathscr{P}^{w}$ to be the root stack of order $w$ along $\infty$ of $\mathbb{P}^{1}$. (See [Cad07, Definition 2.2.4], where this is notated as $\mathbb{P}^{1}_{(\infty,w)}$.) Using [Ols16, Theorem 10.3.10(ii)], the fiber of this root stack over $\infty$ is the stack quotient $\left(\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})\right)/\mu_{r}]$ of the relative spectrum $\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})$ by $\mu_{r}$. Let $\widetilde{\infty}:B\to\mathscr{P}^{w}$ denote the section over $\sigma$ corresponding to map $B\to[\left(\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})\right)/\mu_{r}]$ given by the trivial $\mu_{r}$ torsor over $B$, $\mu_{r}\to B$, and the $\mu_{r}$ equivariant map $\mu_{r}\to B\to\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})$. We use notation as in § 2.1,
@@
where $D,i,X,$ and $h$ satisfy the properties listed in § 2.1. We also assume the order of inertia of $h$ along $\infty$ is $w$ and define $\widetilde{\infty}_{T}$ to be the base change of the section $\widetilde{\infty}$ defined above to $T$. We additionally impose the condition that $h^{\prime}$ is a finite locally free $G$-cover, étale over $\widetilde{\infty}_{T}$, such that the composition of $h^{\prime}:X\to\mathscr{P}^{w}_{T}$ with the coarse space map $\mathscr{P}^{w}_{T}\to\mathbb{P}^{1}_{T}$ is $h$, and $t:T\to X\times_{h^{\prime},\mathscr{P}^{w}_{T},\widetilde{\infty}_{T}}T$ is a section of $h^{\prime}$ over $\widetilde{\infty}_{T}$.
@@
In the definition of pointed Hurwitz space, § 2.1, we choose a particular section $\widetilde{\infty}:B\to\mathscr{P}^{w}$ over $\infty:B\to\mathscr{P}^{1}_{B}$. However, one can show that if we chose a different section $\widetilde{\infty}^{\prime}:B\to\mathscr{P}^{w}$, the resulting pointed Hurwitz space would be isomorphic. The idea is that the sections $\widetilde{\infty}$ and $\widetilde{\infty}^{\prime}$ factor respectively through maps $\alpha:B\to[B/\mu_{w}]$ and $\alpha^{\prime}:B\to[B/\mu_{w}]$, and so it suffices to produce a map $\beta:[B/\mu_{w}]\to[B/\mu_{w}]$ so that $\alpha\circ\beta=\alpha^{\prime}$, as this map will induce a map taking $\widetilde{\infty}$ to $\widetilde{\infty}^{\prime}$, which then induces an isomorphism of these two Hurwitz spaces. The desired map $\beta$ is simply given as the composition of the structure map $B/\mu_{w}]\to B$ with $\alpha^{\prime}$.
@@
We use $\operatorname{CHur}^{G,c}_{n,B}$ to denote the open and closed subscheme of $\operatorname{Hur}^{G,c}_{n,B}$ parameterizing covers $X\to\mathbb{P}^{1}_{T}$ so that the geometric fibers of $X$ over $T$ are connected. We use $\operatorname{CHur}^{G,c}_{n,B}/G]$ to denote the open and closed substack of $\operatorname{Hur}^{G,c}_{n,B}/G]$ parameterizing covers $X\to\mathbb{P}^{1}_{T}$ so that the geometric fibers of $X$ over $T$ are connected.
@@
We will often work over the complex numbers, and hence it will be convenient to assume $B=\operatorname{Spec}\mathbb{C}$. Therefore, we define $\operatorname{Hur}^{G,c}_{n}/G]:=[\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}/G],\operatorname{Hur}^{G,c}_{n}:=\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}},[\operatorname{CHur}^{G,c}_{n}/G]:=[\operatorname{CHur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}/G],\operatorname{CHur}^{G,c}_{n}:=\operatorname{CHur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}$. When $R$ is a ring, we often use use $\operatorname{Hur}^{G,c}_{n,R}$ in place of $\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}R},$ and use analogous notation for other variants of this Hurwitz space.